# Hyp-OW: Exploiting Hierarchical Structure Learning with Hyperbolic Distance Enhances Open World Object Detection

### v1

@@ 4.2 SuperClass Regularization
Let’s denote $\mathcal{S}_{p}$ the set of class indexes belonging to Category $p=1...P$. We approximate the category $p$ embedding by computing the Hyperbolic Average (Khrulkov et al., 2020b) (dubbed HypAve) of every embedding $\{\mathbf{z_{i}^{c}}\}_{i\in\mathcal{M}}$ of classes $c\in\mathcal{S}_{p}$ from the buffer $\mathcal{M}$, that is:
$HypAve(\{\mathbf{z_{i}^{c}}\}_{i\in\mathcal{M},c\in\mathcal{S}_{p}})=\frac{\displaystyle{\sum_{i\in\mathcal{M},c\in\mathcal{S}_{p}}}\gamma_{i}\mathbf{z_{i}^{c}}}{\displaystyle{\sum_{i\in\mathcal{M}}}\gamma_{i}}$ (4)
where $\gamma_{i}=\frac{1}{\sqrt{1-c\lVert x_{i}\rVert^{2}}}$ are the Lorentz factors. To de-clutter the notation, we will denote $\mathbf{\overline{z}_{p}}=HypAve(\{\mathbf{z_{i}^{c}}\}_{i\in\mathcal{M},c\in\mathcal{S}_{p}})$ the Hyperbolic Average feature representation of category $p$. For every element $\mathbf{z_{i}^{c}}$ of a batch $\mathcal{B}$, we sample its category embedding $\mathbf{\overline{z}_{p}}(c\in\mathcal{S}_{p})$ from the buffer $\mathcal{M}$ and use the same contrastive loss:
@@
This loss encourages the features $\mathbf{z}_{i}^{c}$ of each class $c$ to be closer to its corresponding category embedding $\mathbf{\overline{z}}_{p}$, while simultaneously pushing it away from embeddings of other categories $\mathbf{\overline{z}}_{k}$, $k\neq p$.
4.3 Adaptive Relabeling of Unknowns with Hyperbolic Distance
We introduce our Adaptive Relabeling module, which dynamically adapts to the batch statistics to effectively retrieve unknown items (leveraging the learned hyperbolic embedding). Recall for the classification task, we have two types of queries: matched queries with ground-truth labels (found with the Hungarian Algorithm), and unmatched queries that may contain unknown items. The matched queries will establish a threshold criterion to select unknown queries from the unmatched set.
[…]
@@ 4.3 Adaptive Relabeling of Unknowns with Hyperbolic Distance
[…]
@@
5 Experiments
In this section, we begin by describing our experimental setup. We then present comparative results against benchmark baselines, followed by in-depth ablation analysis of each component of Hyp-OW . Due to space limitations, we will defer detailed information to the Appendix Section C and  D.
5.1 Experimental Setup
Datasets
We consider two benchmarks from the literature: the OWOD Split Joseph et al. (2021) and the OWDETR Split Gupta et al. (2022). While the latter (OWDETR Split) strictly separates superclasses across tasks the first (OWOD) has mild semantic overlap between knowns and unknowns across tasks (See Appendix Section A). To validate our hypothesis regarding the semantic relationship between knowns and unknowns, we introduce a third dataset called the Hierarchical Split. This dataset ensures that each task includes at least one class from each category, promoting a higher level of semantic similarity. This will be discussed in paragraph 5.2. Each dataset is defined by four tasks $t=1,2,3,4$, containing 20 labelled classes each, for a total of 80 classes. When task $t$ starts, only the label of classes belonging to that task are revealed. For instance, task $1$ only contains labels of classes from 0 to 19, while task $2$ only contains labels of classes from 21 to 39, and so on.
Implementation Details
We use Deformable DETR (Zhu et al., 2021) pretrained in a self-supervised manner (DINO Caron et al. (2021)) on Resnet-50 (He et al., 2016) as our backbone. The number of deformable transformer encoder and decoder layers are set to $6$. The number of queries is set to $Q=100$ with a dimension $d=256$. During inference time, the top-100 high scoring queries per image are used for evaluation. Additional detail are provided in Appendix Section B.
Metrics and Baselines
Following current metrics used for OWOD, we use mean average precision (mAP) for known items while U-Recall is the main metric used to quantify the unknown detection quality or each method Gupta et al. (2022); Wu et al. (2022a); Zohar et al. (2022); Maaz et al. (2022); Yu et al. (2022). Additional metric is discussed in Table 9 We consider the following baselines from literature: ORE-EBUI (Joseph et al., 2021), UC-OWOD (Wu et al., 2022b), OCPL (Yu et al., 2022), 2B-OCD (Wu et al., 2022a), OW-DETR (Gupta et al., 2022) and PROB (Zohar et al., 2022).
5.2 Benchmark Results
Dataset Similarity
To gain insights into the structure of each dataset, we introduce a semantic similarity measure based on GloVe’s embedding Pennington et al. (2014) (defined in Appendix Section A.1). This measure quantifies the similarity overlap between known and unknown items, with higher values indicating larger overlap. The three splits, OW-DETR Split (Low regime ), OWOD Split (Medium regime ), and Hierarchical Split (High regime ), exhibit a monotonic ranking in terms of similarity, with respective values of $0.27$, $0.33$, and $0.41$. This serves as a starting point for evaluating baseline methods under various scenarios.
Unknown Detection (U-Recall)
Table 1 shows the high performance gain of Hyp-OW over PROB on Medium regime and High regime of $3\%$ on average. This highlights the utility of learning hierarchical structured representations and retrieving unknowns based on their similarity with known objects, as opposed to PROB, which learns a single mean representation for all objects. For the Low regime our method is performing on-par with PROB except for task 1 which shows a surprising improvement of $6$ points. A possible explanation may come from the nature of the Object Detection (OD) task, where there can be significant overlap between bounding boxes. This encourages the model to learn classes that frequently co-occur, such as "person" and "backpack" or "teddy bear" and "bicycle" (See Figure 3). We provide qualitative and quantitative explanations in Appendix Section C.
Task 1 Task 2 Task 3 Task 4
Low
ORE - EBUI 1.5 61.4 3.9 40.6 3.6 33.7 31.8
@@ 5.2 Benchmark Results
PROB 17.6 73.4 (+0.7) 22.3 50.4 24.8 42.0 39.9
Hyp-OW (Ours) 23.9 (+6.3) 72.7 23.3 (+1.0) 50.6 25.4 46.2 (+4.2) 44.8 (+4.9)
@@
Figure 3: Bounding boxes overlap. In the OD task, the high overlap between bounding boxes of frequently co-occurring objects can influence the model to learn their associations and correlations.
@@
From the known accuracy, Hyp-OW surpasses baseline benchmark on all tasks of Hierarchical Split and shows notable performance for the two last tasks of OW-DETR Split. This can be credited to the structural hierarchy learned that groups together class of same category (See t-SNE Figure 1 middle and right).
5.3 Ablation Analysis
We now propose an in-depth understanding of Hyp-OW by removing one by one each component and see its direct impact (illustrated quantitatively in Table 2 for Hierarchical Split).
@@ 5.3 Ablation Analysis
w/ Cosine Distance 32.8 (-2.1) 49.0 (-0.9) 46.4 (-1.1) 45.4 55.4 43.2 (-1.1) 43.1
w/o SuperClass Regularizer 32.0 (-2.9) 50.0 47.1 45.1 52.9 (-2.3) 43.7 43.5
w/o Adaptive Relabeling 34.7 41.2 (-8.7) 47.6 38.9 (-6.6) 54.1 (-1.1) 36.5 (-7.8) 36.1 (-7.8)
Table 2: Impact of each component of Hyp-OW on hierarchical Split. We observe that the Relabeling module (third line) significantly reduces mAP while maintaining U-Recall. On the other hand, the SuperClass Regularizer and Cosine Distance have primarily an impact on unknown detection.
@@
This module uses Eq 6 to relabel unmatched bounding boxes as unknowns. To assess the impact of this relabeling strategy, we compare it with an alternative approach used by PROB Zohar et al. (2022), where all unmatched queries are labeled as unknowns. Although the decrease in U-Recall is marginal, we observe a significant reduction in known accuracy (mAP). This performance degradation can be attributed to the over-prediction of patches as unknowns, which results in misclassification of known objects. Heatmaps in Figure 10 and  11 (Appendix) demonstrate the effectiveness of our module where we observe that unknowns belonging to the same category (sharing some color) as knowns exhibit lower Hyperbolic Distance.
@@
By setting $\beta=0$, we no longer enforce the grouping of items from the same category in the hyperbolic space (compare t-SNE plots in Figure 1). As a result, we observe a reduction in U-Recall of $2.9$, $0.4$, and $2.3$ points, respectively (Table 2, second line). Heatmap Figure 4 illustrates the hyperbolic distance from each class to every category’s embedding (computed using Eq 4) with lighter colors indicating smaller distances. With our Regularizer (right plot), we observe a wider range of values spanning from $0.7$ to $2.30$, compared to a smaller range of $0.78$ to $1.2$ without the Regularizer. This highlights the effect of our Regularizer that pushes classes from different categories apart (shown as dark color in the right plot) while bringing classes of similar categories closer together. A more detailed plot can be found in Appendix Figure 12.
Cosine Distance:
Trading hyperbolic distance for the cosine distance ($c=0$) harms both the U-Recall and mAP as the hyperbolic embedding space is more suitable to learn data with latent hierarchical structure (highlighted by Figure 14 in Appendix).
6 Conclusion
@@
The Open World Object Detection framework presents a challenging and promising setting, encompassing crucial aspects such as lifelong learning and unknown detection. In our work, we have emphasized the lack of a clear definition of unknowns and the need of a hierarchical or semantic relationship between known and unknown classes. This led us to propose Hyp-OW that focuses on learning and modeling the structural hierarchy within the dataset, which is then utilized for unknowns retrieval. Extensive experiments demonstrates significant improvement of Hyp-OW for both known and unknown detection (up to 6 points) particularly in the presence of inherent hierarchy between classes.
Acknowledgement
We would like to thank (in no particular order): Jiajing Guo, Shabnam Ghaffarzadegan, Yulian Guo, Sharath Gopal, Sanbao Bu and Jorge Ono for useful discussions and feedbacks throughout the work.
Appendix
The Appendix Section is organized as follow:
@@ Acknowledgement
  • Section A provides information on the composition of each dataset and the introduction of our semantic similarity measure.
  • Section B presents the experimental details.
@@
  • Section C includes comprehensive benchmark metrics results and visual plots (heatmaps) to illustrate our findings.
  • Section D discusses the results of the ablation analysis accompanied by visual figures for better understanding.
Appendix A Dataset Information
A.1 Dataset Composition

### v2

@@ Introduction
Let’s denote $\mathcal{S}_{p}$ as the set of class indexes within categories $p=1...P$ (we denote this set $\mathcal{P}$). For instance, the category ”vehicles” might encompass classes such as car, truck, bus, and so on… We approximate the category $p$ embedding by computing the Hyperbolic Average (Khrulkov et al. 2020b) (dubbed HypAve) of every embedding $\{\mathbf{z_{i}^{c}}\}_{i\in\mathcal{M}}$ of classes $c$ from the buffer $\mathcal{M}$ within this category ($c\in\mathcal{S}_{p}$) that is:
$\mathbf{\overline{z}_{p}}=HypAve(\{\mathbf{z_{i}^{c}}\}_{i\in\mathcal{M},c\in\mathcal{S}_{p}})=\frac{\displaystyle{\sum_{i\in\mathcal{M},c\in\mathcal{S}_{p}}}\gamma_{i}\mathbf{z_{i}^{c}}}{\displaystyle{\sum_{i\in\mathcal{M}}}\gamma_{i}}$ (4)
where $\gamma_{i}=\frac{1}{\sqrt{1-c\lVert x_{i}\rVert^{2}}}$ is the Lorentz factor. For each element $\mathbf{z_{i}^{c}}$ of a batch $\mathcal{B}$, we extract its category embedding $\mathbf{\overline{z}_{p}}$ $(c\in\mathcal{S}_{p})$ from the buffer $\mathcal{M}$. Using a temperature parameter $\tau_{2}$, we formulate our SuperClass regularizer as follows:
@@
This loss encourages the features $\mathbf{z}_{i}^{c}$ of each class $c$ to be closer to its corresponding category embedding $\mathbf{\overline{z}}_{p}$, while simultaneously pushing it away from embeddings of other categories $\mathbf{\overline{z}}_{k}$, $k\neq p$. In essence, it fosters the grouping of similar items at the SuperClass/category level rather than the individual class level.
Adaptive Relabeling of Unknowns with Hyperbolic Distance
We introduce our Adaptive Relabeling module, which dynamically adapts to the batch statistics to effectively detect unknowns. We can summarize this procedure in three steps: a) the hyperbolic mean, or centroid, is calculated for each class in $\mathcal{M}$, b) for all known items in an image, we compute the distance to each centroid, the greatest distance is labeled as $\delta_{B}$, c) for every unmatched bounding box, we calculate its distance to each centroid: if less than $\delta_{B}$, it is relabeled as ’unknown’ otherwise it is considered as a background.
a) We define $\mathbf{\underline{z}_{c}}$  [^3] We differentiate from $\mathbf{\overline{z}_{p}}$ with an underline to distinguish Hyperbolic Average of class and category the hyperbolic average of class $c$ computed from the buffer $\mathcal{M}$ as:
@@
b) We now use the matched queries to define: $\delta_{\mathcal{B}}=\displaystyle{\max_{i\in\mathcal{B},c\in\mathcal{K}}}\;d_{hyp}(\mathbf{z_{i}},\mathbf{\underline{z}_{c}})$. In essence, $\delta_{\mathcal{B}}$ signifies the greatest distance from any known items in the batch $\mathcal{B}$ to all centroid $\mathbf{\underline{z}_{c}},c\in\mathcal{K}$ in the replay buffer $\mathcal{M}$.
c) This threshold is then utilized to relabel any unmatched query $\mathbf{z_{u}}$ as unknown if:
$\displaystyle{\min_{c\in\mathcal{K}}}\;d_{hyp}(\mathbf{z_{u}},\mathbf{\underline{z}_{c}})\leq\delta_{\mathcal{B}}$ (6)
@@
Experiments
In this section, we start with describing our experimental setup. We then present comparative results against benchmark baselines, followed by in-depth ablation analysis of each component of Hyp-OW . Due to space limitations, we will defer detailed information to the Supplementary Material.
Experimental Setup
Regime Task 1 Task 2 Task 3 Task 4
Methods U-Recall ($\uparrow$) mAP ($\uparrow$) U-Recall ($\uparrow$) mAP ($\uparrow$) U-Recall ($\uparrow$) mAP ($\uparrow$) mAP ($\uparrow$)
Low ORE - EBUI 1.5 61.4 3.9 40.6 3.6 33.7 31.8
@@
PROB 17.6 73.4 22.3 50.4 24.8 42.0 39.9
Hyp-OW (Ours) 23.9 72.7 23.3 50.6 25.4 46.2 44.8
$\Delta$(Rel. Difference) +6.3 $\leq 1.0$ +1.0 $\leq 1.0$ $\leq 1.0$ +4.2 +4.9
@@
Implementation Details
We use Deformable DETR (Zhu et al. 2021) pretrained in a self-supervised manner (DINO (Caron et al. 2021)) on Resnet-50 (He et al. 2016) as our backbone. The number of deformable transformer encoder and decoder layers are set to $6$. The number of queries is set to $Q=100$ with a dimension $d=256$. During inference time, the top-100 high scoring queries per image are used for evaluation. For our method, We used $c=0.1$, $\tau_{1}=0.2$, $\tau_{2}=0.4$. For the set $S_{p}$ that defines the composition of each category (SuperClass), we adhere to the grouping used in MS-COCO dataset (Lin et al. 2014). All used hyperparameters can be found in the Supplementary.
Metrics and Baselines
Following the current metrics used for OWOD, we utilize the mean average precision (mAP) for known items, while U-Recall serves as the primary metric to quantify the quality of unknown detection for each method (Gupta et al. 2022; Wu et al. 2022a; Zohar, Wang, and Yeung 2023; Maaz et al. 2022; Yu et al. 2022). Additional metric is discussed in Table 4. We consider the following baselines from literature: OW-DETR (Gupta et al. 2022) and PROB (Zohar, Wang, and Yeung 2023). While we included Faster R-CNN methods as informative references (ORE-EBUI (Joseph et al. 2021), UC-OWOD (Wu et al. 2022b), OCPL (Yu et al. 2022), 2B-OCD (Wu et al. 2022a)), our primary emphasis is on comparing against deformable Transformer-based methods to ensure a fair assessment following the evaluation procedure of PROB.
Datasets
We consider two benchmarks from the literature: the OWOD Split (Joseph et al. 2021) and the OWDETR Split (Gupta et al. 2022). While the latter (OWDETR Split) strictly separates SuperClasses across tasks the first (OWOD) has mild semantic overlap between knowns and unknowns across tasks (See Supplementary Material). To closely mimic real-world scenarios, we consider a Hierarchical Split which ensures that each task includes at least one class from each category [^4] While there are various ways to distribute the classes across tasks, our primary intention here is to introduce a third scenario for assessing all methods.. Each dataset is defined by four tasks $t=1,2,3,4$, containing 20 labelled classes each, for a total of 80 classes. When task $t$ starts, only the label of classes belonging to that task are revealed. For instance, task $1$ only contains labels of classes from 0 to 19, while task $2$ only contains labels of classes from 21 to 39, and so on. Composition of each dataset can be found in the Supplementary.
Dataset Structure
To better understand the structure of each dataset, we define a semantic similarity measure using GloVe’s embedding (Pennington, Socher, and Manning 2014). Denoting $\mathbf{\omega_{c}},c\in\mathcal{K}$ ( respectively $\mathbf{\omega_{k}},c\in\mathcal{U}$ ) the embedding of known (unknown) classes, the semantic overlap between knowns and unknowns for task $t\in[1,T-1]$ is:
$S_{t}=\frac{1}{|\mathcal{U}^{t}|}\displaystyle{\sum_{k\in\mathcal{U}^{t}}\max_{c\in\mathcal{K}^{t}}}\frac{<\mathbf{\omega^{c}},\mathbf{\omega^{k}>}}{\lVert\mathbf{\omega^{c}}\rVert\cdot\lVert\mathbf{\omega^{k}}\rVert}$ (8)
Higher values indicate larger overlap between the knowns and unknowns. Figure 3 effectively quantifies the level of hierarchical structure inherent in each dataset, providing a clear insight into the composition of each split. This framework now offers us a consistent basis to evaluate each method across different hierarchical scenarios: Low regime (OW-DETR Split), Medium regime (OWOD Split) and High regime (Hierarchical Split). This metric is increasing as the number of knowns grows throughout the training
Benchmark Results
Unknown Detection (U-Recall)
Table 1 shows the high performance gain of Hyp-OW over PROB on Medium regime and High regime of $3\%$ on average (The row $\Delta$ indicates relative performance with respect to the second best algorithm). This highlights the utility of learning hierarchical structured representations and retrieving unknowns based on their similarity with known objects, as opposed to PROB, which learns a single mean representation for all objects. For the Low regime our method is performing on-par with PROB except for task 1 which shows a surprising improvement of $6$ points. Overall, Hyp-OW demonstrates consistent and strong results on the three benchamrks.
@@
Hyp-OW outperforms baseline benchmarks across all tasks in the Hierarchical Split and notably enhances performance for the last two tasks in the OW-DETR Split. This success can be attributed to the learned structural hierarchy, which groups classes of the same category (illustrated in t-SNE Figure 1). Moreover, our method exhibits robust performance even in the low regime. This can be attributed to the inherent presence of bounding box overlaps in object detection tasks, enabling the model to learn about co-occurring objects (refer to Supplementary Material).
U-Recall ($\uparrow$) mAP($\uparrow$)
$c=0.0$ (Cosine Dist.) 32.8 49.0
$c=0.1$ (Hyp-OW ) 34.9 49.9
$c=0.2$ 33.3 49.5
$c=0.5$ 32.3 49.8
Table 2: Impact of curvature coefficient $c$ for Hierarchical Split Task 1($\uparrow$ indicates larger values are better and $\downarrow$ indicates smaller values are better).
Figure 4: t-SNE plot of the learned class representations Hyperbolic Distance tends to learns a better hierarchical structure than Cosine Distance.
@@
w/ Cosine Distance (c=0) 32.8 49.0 46.4 45.4 55.4 43.2 43.1
w/o SuperClass Regularizer 32.0 50.0 47.1 45.1 52.9 43.7 43.5
w/o Adaptive Relabeling 34.7 41.2 47.6 38.9 54.1 36.5 l 36.1
Table 3: Impact of each component of Hyp-OW on Hierarchical Split. We observe that the Relabeling module (third line) significantly reduces mAP while maintaining U-Recall. On the other hand, the SuperClass Regularizer and Cosine Distance have primarily an impact on unknown detection. Task 4 does not have U-Recall since all 80 classes are known at this stage.
Ablation Analysis
We now aim to gain an in-depth understanding of Hyp-OW by systematically removing each component one by one to assess its direct impact (illustrated quantitatively in Table 3 for Hierarchical Split and in the Supplementary for OWOD-Split). Additionally, we perform a quantitative analysis of the unknown confusion. Qualitative visualizations can be found in the supplementary.
Curvature Coefficient $c$:
We assess the impact of different hyperbolic distances alongside the cosine distance $c=0$ (which is also linked with the Euclidian distance for normalized vectors, See Supplementary) in Table 2. While substituting hyperbolic distance with cosine distance ($c=0$) negatively affects both U-Recall and mAP, higher $c$ degrades mainly the U-Recall. The hyperbolic embedding space is more suitable to learn data with latent hierarchical structure (see t-SNE plot Figure 4).
@@
The relabeling of unmatched bounding boxes as unknowns is governed by Eq 6. To evaluate its impact, we constrast it with an alternative technique used by PROB (Zohar, Wang, and Yeung 2023), where all unmatched queries are classified as unknowns. Results are shown in Table 3 fourth row. Although the decrease in U-Recall is marginal, a noteworthy reduction in known accuracy (mAP) is observed. This decline can be attributed to the over-prediction of patches as unknowns, which results in misclassification of known objects. More detailed figures in the Supplementary show its efficacy where we notice that unknowns belonging to the same category as knowns exhibit lower Hyperbolic Distance, manifested as lighter colors.
@@
By setting $\beta=0$ (Table 3: third row), we no longer enforce the grouping of items at the category level (compare t-SNE plots in Figure 1). We then observe a reduction in U-Recall of $2.9$, $0.4$, and $2.3$ points respectively. Heatmap Figure 5 illustrates the hyperbolic distance from each class to every category’s embedding (computed using Eq 4) with lighter colors indicating smaller distances. With our regularizer (right plot), we observe a wider range of values spanning from $0.7$ to $2.30$, compared to a smaller range of $0.78$ to $1.2$ without the regularizer. This emphasizes the impact of our regularizer, which effectively separates classes from distinct categories (depicted by the darker color in the right plot) while simultaneously bringing classes from similar category closer together. A more detailed plot can be found in the Supplementary.
Unknowns Confusion:
We measure the A-OSE metric introduced by  Joseph et al. (2021) (defined in the Supplementary Material) which quantifies the number of unknowns misclassified as knowns (lower is better). Table 4 showcases the results for Hierarchichal Split (see in Supplementary Material for OWOD Split). Comparing with PROB, Hyp-OW exhibits significantly fewer misclassifications across different tasks, with at least a twofold reduction for the first two tasks and approximately $20\%$ less for the third task.
Task 1 Task 2 Task 3
A-OSE($\downarrow$) A-OSE($\downarrow$) A-OSE($\downarrow$)
OW-DETR 42,540 26,527 20,034
PROB 14,962 8,929 5,387
Hyp-OW (Ours) 7,420 3,849 4,611
Table 4: A-OSE metric on Hierarchical Split. Hyp-OW exhibits a lower rate of unknowns misclassifition as knowns compared to other baselines.
Conclusion
@@
The Open World Object Detection framework presents a challenging and promising setting, encompassing crucial aspects such as lifelong learning and unknown detection. In our work, we have emphasized the lack of a clear definition of unknowns and the need for a hierarchical or semantic relationship between known and unknown classes. This led us to propose Hyp-OW that focuses on learning and modeling the structural hierarchy within the dataset, which is then utilized for unknowns retrieval. Extensive experiments demonstrate significant improvement of Hyp-OW for both known and unknown detection (up to 6 percent) particularly in the presence of inherent hierarchy between classes. Future directions include leveraging knowledge from pretrained vision language models to detect desired unknowns. We also hope that our hierarchical structural learning paradigm benefits adjacent fields such as OOD detection (Behpour et al. 2023), data pre-selection (Li et al. 2023) or instance segmentation (Wang et al. 2022).
Appendix A Supplementary Material
In the next secions we respectively cover the following topics:
@@ Appendix A Supplementary Material
  • Dataset composition for each split and introduction of our semantic similarity measure
  • Experimental details
  • Qualitative visualization demonstrating the effectiveness of Hyp-OW in comparison to PROB and individual components of Hyp-OW
@@
  • Comprehensive benchmark metrics results and visual plots (heatmaps) to illustrate our findings
  • Ablation analysis discussion accompanied with their qualitative plots
Appendix B Dataset Information
Dataset Composition